If Tn is the nth term of an exponential sequence, show that T1Tn=T2Tn-1
Answer by ikleyn(52754)

.

Let T be the first term of the sequence and  q be the common ratio.



We have  T1 = T;  Tn = Tq^(n-1);  T2 = Tq  and  T(n-1) = TQ^(n-2).



Substitute these values into the left side and into the right side, and you will get the same  
expression T^2*q^(n-1) in both sides.